# Randomly punctured Reed–Solomon codes achieve list-decoding capacity over linear-sized fields

### v5

@@ Abstract
A recent breakthrough by Brakensiek, Gopi, and Makam, established that Reed–Solomon codes are combinatorially list-decodable all the way to capacity. However, their results hold for randomly-punctured Reed–Solomon codes over an exponentially large field size $2^{O(n)}$, where $n$ is the block length of the code. A natural question is whether Reed–Solomon codes can still achieve capacity over smaller fields. Recently, Guo and Zhang showed that Reed–Solomon codes are list-decodable to capacity with field size $O(n^{2})$. We show that Reed–Solomon codes are list-decodable to capacity with linear field size $O(n)$, which is optimal up to the constant factor. We also give evidence that the ratio between the alphabet size $q$ and code length $n$ cannot be bounded by an absolute constant.
Our techniques also show that random linear codes are list-decodable up to (the alphabet-independent) capacity with optimal list-size $O(1/\varepsilon)$ and near-optimal alphabet size $2^{O(1/\varepsilon^{2})}$, where $\varepsilon$ is the gap to capacity. As far as we are aware, list-decoding up to capacity with optimal list-size $O(1/\varepsilon)$ was not known to be achievable with any linear code over a constant alphabet size (even non-constructively), and it was also not known to be achievable for random linear codes over any alphabet size.
Our proofs are based on the ideas of Guo and Zhang, and we additionally exploit symmetries of reduced intersection matrices. With our proof, which maintains a hypergraph perspective of the list-decoding problem, we include an alternate presentation of ideas from Brakensiek, Gopi, and Makam that more directly connects the list-decoding problem to the GM-MDS theorem via a hypergraph orientation theorem.
@@ 1 Introduction
An (error-correcting) code is simply a set of strings (codewords). In this paper, all codes are linear, meaning our code $C\subset\mathbb{F}_{q}^{n}$ is a space of vectors over a finite field $\mathbb{F}_{q}$, for some prime power $q$. A Reed–Solomon code [RS60] is a linear code obtained by evaluating low-degree polynomials over $\mathbb{F}_{q}$. More formally,
@@
The rate $R$ of a code $C$ is $R\stackrel{{\scriptstyle\rm def}}{{=}}\log_{q}|C|/n$, which, for a Reed–Solomon code, is $k/n$. Famously, Reed–Solomon codes are optimal for the unique decoding problem [RS60]: for any rate $R$ Reed–Solomon code, for every received word $y\in\mathbb{F}_{q}^{n}$, there is at most one codeword within Hamming distance $pn$ of $y$, [^1] The Hamming distance between two codewords is the number of coordinates on which they differ. and this error parameter $p=\frac{1-R}{2}$ is optimal by the Singleton bound [Sin64].
@@
In this paper, we study Reed–Solomon codes in the context of list-decoding, a generalization of unique-decoding that was introduced by Elias and Wozencraft [Eli57, Woz58]. Formally, a code $C\subset\mathbb{F}_{q}^{n}$ is $(p,L)$-list-decodable if, for every received word $y\in\mathbb{F}_{q}^{n}$, there are at most $L$ codewords of $C$ within Hamming distance $pn$ of $y$.
@@
It is well known that the list-decoding capacity, namely the largest fraction of errors that can be list-decoded with small lists, is $1-R$ [GRS22, Theorem 7.4.1]. Specifically, for $p=1-R-\varepsilon$, there are (infinite families of) rate $R$ codes that are $(p,L)$ list-decodable for a list-size $L$ as small as $O(1/\varepsilon)$. On the other hand, for $p=1-R+\varepsilon$, if a rate $R$ code is $(p,L)$ list decodable, the list size $L$ must be exponential in the code length $n$. Informally, a code that is list-decodable up to radius $p=1-R-\varepsilon$ with list size $O_{\varepsilon}(1)$, or even list size $n^{O_{\varepsilon}(1)}$ where $n$ is the code length, is said to achieve (list-decoding) capacity.
@@
The list-decodability of Reed–Solomon codes is important for several reasons. Reed–Solomon codes are the most fundamental algebraic error-correcting code. In fact, all of the prior explicit constructions of codes achieving list-decoding capacity are based on algebraic constructions that generalize Reed–Solomon codes, for example, Folded Reed–Solomon codes [GR08, KRZSW18], Multiplicity codes [GW13, Kop15, KRZSW18], and algebraic-geometric codes [DL12, GX12, GX13, HRZW19]. Thus, it is natural to wonder whether and when Reed–Solomon codes themselves achieve list-decoding capacity. Additionally, all Reed–Solomon codes are optimally unique-decodable, so (equivalently) they are optimally list-decodable $L=1$, making them a natural candidate for codes achieving list-decoding capacity. Further, capacity-achieving Reed–Solomon codes would potentially offer advantages over existing explicit capacity-achieving codes, such as simplicity and potentially smaller alphabet sizes (which we achieve in this work). Lastly, list-decoding of Reed–Solomon codes has found several applications in complexity theory and pseudorandomness [CPS99, STV01, LP20].
@@
For all these reasons, the list-decodability of Reed–Solomon codes is well-studied. As rate $R$ Reed–Solomon codes are uniquely decodable up to the optimal radius $\frac{1-R}{2}$ given by the Singleton Bound, the Johnson-bound [Joh62] automatically implies that Reed–Solomon codes are $(p,L)$-list-decodable for error parameter $p=1-\sqrt{R}-\varepsilon$ and list size $L=O(1/\varepsilon)$. Guruswami and Sudan [GS99] showed how to efficiently list-decode Reed–Solomon codes up to the Johnson radius $1-\sqrt{R}$. For a long time, this remained the best list-decodability result (even non-constructively) for Reed–Solomon codes.
@@
Since then, several results suggested Reed–Solomon codes could not be list-decoded up to capacity, and in fact, not much beyond the Johnson radius $1-\sqrt{R}$. Guruswami and Rudra [GR06] showed that, for a generalization of list-decoding called list-recovery, Reed–Solomon codes are not list-recoverable beyond the (list-recovery) Johnson bound in some parameter settings. Cheng and Wan [CW07] showed that efficient list-decoding of Reed–Solomon codes beyond the Johnson radius in certain parameter settings implies fast algorithms for the discrete logarithm problem. Ben-Sasson, Kopparty, and Radhakrishnan [BKR10] showed that full-length Reed–Solomon codes ($q=n$) are not list-decodable much beyond the Johnson bound in some parameter settings.
@@
Since then, an exciting line of work [RW14, ST20, ${}^{+}$22, FKS22, GST22, BGM23, GZ23] has shown the existence of Reed–Solomon codes that could in fact be list-decoded beyond the Johnson radius. These works all consider combinatorial list-decodability of randomly punctured Reed–Solomon codes. By combinatorial list-decodability, we mean that the code is proved to be list-decodable without providing an algorithm to efficiently decode the list of nearby codewords. By randomly punctured Reed–Solomon code, we mean a code $\mathsf{RS}_{n,k}(\alpha_{1},\dots,\alpha_{n})$ where $(\alpha_{1},\dots,\alpha_{n})$ are chosen uniformly over all $n$-tuples of pairwise distinct elements of $\mathbb{F}_{q}$. Several of these works [RW14, FKS22, GST22] proved more general list-decoding results about randomly puncturing any code with good unique-decoding properties, not just Reed–Solomon codes.
@@
In this line of work, a recent breakthrough of Brakensiek, Gopi, and Makam [BGM23] showed, using notions of “higher-order MDS codes” [BGM22, Rot22], that Reed–Solomon codes can actually be list-decoded up to capacity. In fact, they show, more strongly, that Reed–Solomon codes can be list-decoded with list size $L$ with radius $p=\frac{L}{L+1}(1-R)$, exactly meeting the generalized Singleton bound [ST20], resolving a conjecture of Shangguan and Tamo [ST20]. However, their results require randomly puncturing Reed–Solomon codes over an exponentially large field size $2^{O(n)}$, where $n$ is the block length of the code.
@@
A natural question is how small we can take the field size in a capacity-achieving Reed–Solomon code. Brakensiek, Dhar, and Gopi [BDG22, Corollary 1.7, Theorem 1.8] showed that the exponential-in-$n$ field size in [BGM23] is indeed necessary to exactly achieve the generalized Singleton bound for $L=2$ — under the additional assumptions that the code is linear and MDS. These assumptions were removed in followup work [AGL24], which also generalized the result to all $L$ — but smaller field sizes remained possible if one allowed a small $\varepsilon$ slack in the parameters. Recently, an exciting work of Guo and Zhang [GZ23] showed that Reed–Solomon codes are list-decodable up to capacity, in fact up to (but not exactly at) the generalized Singleton bound, with alphabet size $O(n^{2})$.
@@ 1.1 Our results
1.1.0.0.1 List-decoding Reed–Solomon codes.
Building on Guo and Zhang’s argument, we show that Reed–Solomon codes are list-decodable up to capacity and the generalized Singleton bound with linear alphabet size $O(n)$, which is evidently optimal up to the constant factor. Our main result is the following.
@@
As in previous works like [BGM23, GZ23], Theorem 1.1 gives average-radius list-decodability, a stronger guarantee than list-decodability: for any distinct $L+1$ codewords $c^{(1)},\dots,c^{(L+1)}$ and any vector $y\in\mathbb{F}_{q}^{n}$, the average Hamming distance from $c^{(1)},\dots,c^{(L+1)}$ to $y$ is at least $\frac{L}{L+1}(1-R-\varepsilon)$. Taking $L=O(1/\epsilon)$ in Theorem 1.1, it follows that Reed–Solomon codes achieve list-decoding capacity even over linear-sized alphabets.
@@
Let $\varepsilon\in(0,1)$ and $q$ be a prime power such that $q\geq n+k\cdot 2^{O(1/\varepsilon^{2})}$. Then with probability at least $1-2^{-\Omega(n/\varepsilon)}$, a randomly punctured Reed–Solomon code of block length $n$ and rate $k/n$ over $\mathbb{F}_{q}$ is $(1-R-\varepsilon,O(\frac{1}{\varepsilon}))$ average-radius list-decodable.
The alphabet size in [GZ23] is $2^{O(L^{2}/\varepsilon)}nk$. Our main contribution is improving their alphabet size from quadratic to linear. As a secondary improvement, we also bring down the constant factor from $2^{O(L^{2}/\varepsilon)}$ to $2^{O(L/\varepsilon)}$. We defer the proof overview of Theorem 1.1 to Section 3.1 after setting up the necessary notions in Section 2.
@@
In our proof of Theorem 1.1, we maintain a hypergraph perspective of the list-decoding problem, which was introduced in [${}^{+}$22]. Section 2.2 elaborates on the advantages of this perspective, which include (i) more conpact notations, definitions, and lemma statements, (ii) our improved constant factor of $2^{O(L/\varepsilon)}$, (iii) an improved alphabet size in our random linear codes result below (Theorem 1.3), and (iv) an alternate presentation of ideas from Brakensiek, Gopi, and Makam [BGM23] that more directly connects the list-decoding problem to the GM-MDS theorem [DSY14, Lov18, YH19] via a hypergraph orientation theorem (see Appendix A).
1.1.0.0.2 List-decoding random linear codes.
@@
A random linear code of rate $R$ and length $n$ over $\mathbb{F}_{q}$ is a random subspace of $\mathbb{F}_{q}^{n}$ of dimension $Rn$. List-decoding random linear codes is well-studied [ZP81, Eli91, GHSZ02, GHK11, Woo13, RW14, RW18, LW20, ${}^{+}$20, ${}^{+}$21, GM22, PP23] and is an important question for several reasons. First, finding explicit codes approaching list-decoding capacity is a major challenge, and random linear codes provide a stepping stone towards explicit codes: a classic result says that uniformly random codes achieve list-decoding capacity [Eli57, Woz58], and showing list-decodability of random linear codes can be viewed as a derandomization of the uniformly random construction. Mathematically, the list-decodability of random linear codes concerns a fundamental geometric question: to what extent do random subspaces over $\mathbb{F}_{q}$ behave like uniformly random sets? In coding theory, list-decodable random linear codes are useful building blocks in other coding theory constructions [GI01, HW18]. Lastly, the algorithmic question of decoding random linear codes is closely related to the Learning With Errors (LWE) problem in cryptography [Reg09] and Learning Parity with Noise (LPN) problem in learning theory [BKW03, FGKP06].
@@
The list-decodability of random linear codes is more difficult to analyze than uniformly random codes, because codewords do not enjoy the same independence as in random codes. Thus the naive argument that shows that random linear codes achieve list-decoding capacity [ZP81] gives an exponentially worse list size of $q^{1/\varepsilon}$ than for random codes ($\varepsilon$ is the gap to the “$q$-ary capacity”, $R=1-H_{q}(p)$, where $H_{q}(x)\stackrel{{\scriptstyle\rm def}}{{=}}x\log_{q}(q-1)-x\log_{q}(x)-(1-x)\log_{q}(1-x)$ is the $q$-ary entropy function). Several works have sought to circumvent this difficulty [Eli91, GHSZ02, GHK11, Woo13, RW14, RW18, LW20, ${}^{+}$21] improving the list-size bound to $O_{q}(1/\varepsilon)$, matching the list-size of uniformly random codes.
@@
However, these results are more relevant for smaller alphabet sizes $q$, and approaching the alphabet-independent capacity of $p=1-R$ is less understood. In this setting, uniformly random codes are, with high probability, list-decodable to capacity with optimal alphabet size $2^{O(1/\varepsilon)}$  [^2] This follows from the list-decoding capacity theorem [Eli57, Woz58]. Over $q$-ary alphabets, the list-decoding capacity is given by $p=H_{q}^{-1}(1-R)$, which is larger than $1-R-\varepsilon$ when $q\geq 2^{\Omega(1/\varepsilon)}$. and optimal list size $O(1/\varepsilon)$. [^3] For codes over smaller alphabets, the list size $O(1/\varepsilon)$, where $\varepsilon$ is the gap to capacity, is believed to be optimal, but a proof is only known for large radius [GV10]. However, for approaching the alphabet independent capacity, the list size $O(1/\varepsilon)$ is known to be optimal by the generalized Singleton bound [ST20]. However, it was not known whether random linear codes (or, in general, more structured codes) could achieve similar parameters. In particular, both of the following questions were open (as far as we are aware).
@@
  • Are rate $R$ random linear codes $(1-R-\varepsilon,O(1/\varepsilon))$-list-decodable with high probability? Previously, this was not known for any alphabet size $q$, even alphabet size growing with the length of the code. Previously, the best list size for random linear codes list-decodable to radius $p=1-R-\varepsilon$ was at least $2^{\Omega(1/\varepsilon)}$ [GHK11, RW18].444[GHK11] appears to give a list-size bound of $O(q^{O_{R}(1)}/\varepsilon)$, and [RW18] appears to give a list size bound that is at least $q^{\log^{2}(1/\varepsilon)}$, and we need $q\geq 2^{\Omega(1/\varepsilon)}$
@@
By taking $L=O(1/\varepsilon)$, we see that random linear codes achieve capacity with optimal list size $O(1/\varepsilon)$ and near optimal alphabet size $2^{O(1/\varepsilon^{2})}$.
@@
The techniques developed in this work for the proof of Theorem 1.1 are important for obtaining the strong alphabet size guarantees of Theorem 1.3. One could also have adapted the proof of Guo and Zhang, but doing so in the same natural way would only yield an alphabet size of $O(n)$ (see Section 4.4 for discussions). Further, our use of the hypergraph machinery, which gives a secondary improvement over [GZ23] in constant factor in the alphabet size in Corollary 1.2, gives the primary improvement in the alphabet size in Corollary 1.4 from $2^{O(1/\varepsilon^{3})}$ to $2^{O(1/\varepsilon^{2})}$.
@@
1.1.0.0.3 Alphabet size lower bounds.
Above, we saw that random linear codes achieve list-decoding capacity with optimal list-size and near-optimal alphabet size. A natural question, asked by Guo and Zhang, is how large the alphabet size needs to be for capacity-achieving Reed–Solomon codes. We showed that $q\geq n\cdot 2^{O(1/\varepsilon^{2})}$ suffices, and by the list-decoding capacity theorem [Eli57, Woz58], we cannot have better than an exponential-type dependence on $1/\varepsilon$ for subconstant $\varepsilon<O(1/\log n)$.
For approaching capacity with constant $\varepsilon$, Ben-Sasson, Kopparty, and Radhakrishnan [BKR10] showed that, for any $c\geq 1$, there exist full-length Reed–Solomon codes that are not list-decodable much beyond the Johnson bound with list-sizes $O(n^{c})$. Thus in order to achieve list-decoding capacity, one needs $q>n$ in some cases. However, while full-length Reed–Solomon codes could not achieve capacity, perhaps it was possible that Reed–Solomon codes over field size, say $q=2n$ or even $q=(1+\gamma)n$, could achieve capacity in all parameter settings. We observe that, as a corollary of [BKR10], such a strong guarantee is not possible. We show that, for any $c>1$, there exist a constant rate $R=R(c)>0$ and infinitely many field sizes $q$ such that all Reed–Solomon codes of length $n\geq q/c$ and rate $R$ over $\mathbb{F}_{q}$ are not list-decodable to capacity $1-R$ with list size $n^{c}$. The proof is in Appendix B.
@@
Follow up work
The techniques in our paper have already been influential. In follow-up work, Brakensiek, Dhar, Gopi, and Zhang [${}^{+}$23b] used our argument to prove that Algebraic Geometry (AG) codes achieve list-decoding capacity over constant-sized alphbaets. They prove this by combining our techniques with a generalized GM-MDS theorem, proved by Brakensiek, Dhar, Gopi [BDG23a].

### v6

@@ Abstract
A recent breakthrough by Brakensiek, Gopi, and Makam established that Reed–Solomon codes are combinatorially list-decodable all the way to capacity. However, their results hold for randomly-punctured Reed–Solomon codes over an exponentially large field size $2^{O(n)}$, where $n$ is the block length of the code. A natural question is whether Reed–Solomon codes can still achieve capacity over smaller fields. We show that Reed–Solomon codes are list-decodable to capacity with linear field size $O(n)$, which is evidently optimal up to a constant factor.
Our techniques also show that random linear codes are list-decodable up to capacity with optimal list-size $O(1/\varepsilon)$ and near-optimal alphabet size $2^{O(1/\varepsilon^{2})}$, where $\varepsilon$ is the gap to capacity. As far as we are aware, list-decoding up to capacity with optimal list-size $O(1/\varepsilon)$ was not known to be achievable with any linear code over a constant alphabet size (even non-constructively), and it was also not known to be achievable for random linear codes over any alphabet size.
With our proof, which maintains a hypergraph perspective of the list-decoding problem, we include an alternate presentation of ideas from Brakensiek, Gopi, and Makam that more directly connects the list-decoding problem to the GM-MDS theorem via a hypergraph orientation theorem.
[…]
[classification=text]
@@ 1 Introduction
An (error-correcting) code is simply a set of strings (codewords). In this paper, all codes are linear, meaning our code $C\subseteq\mathbb{F}_{q}^{n}$ is a space of vectors over a finite field $\mathbb{F}_{q}$, for some prime power $q$. A Reed–Solomon code [54] is a linear code obtained by evaluating low-degree polynomials over $\mathbb{F}_{q}$. More formally,
@@
The rate $R$ of a code $C$ is $R\stackrel{{\scriptstyle\rm def}}{{=}}\log_{q}|C|/n$, which, for a Reed–Solomon code, is $k/n$. Famously, Reed–Solomon codes are optimal for the unique decoding problem [54]: for any rate $R$ Reed–Solomon code, for every received word $y\in\mathbb{F}_{q}^{n}$, there is at most one codeword within Hamming distance $pn$ of $y$ for error parameter p=(1R)/2p=(1-R)/2, [^1] The Hamming distance between two codewords is the number of coordinates on which they differ. and further this error parameter $p=\frac{1-R}{2}$ is optimal by the Singleton bound [60].
@@
In this paper, we study Reed–Solomon codes in the context of list-decoding, a generalization of unique-decoding that was introduced by Elias and Wozencraft [16, 64]. Formally, a code $C\subseteq\mathbb{F}_{q}^{n}$ is $(p,L)$-list-decodable if, for every received word $y\in\mathbb{F}_{q}^{n}$, there are at most $L$ codewords of $C$ within Hamming distance $pn$ of $y$.
@@
It is well known that the largest fraction of errors that can be list-decoded with small lists approaches the quantity 1R1-R [35, Theorem 7.4.1]. Specifically, for $p=1-R-\varepsilon$, there are (infinite families of) rate $R$ codes that are $(p,L)$ list-decodable for a list-size $L$ as small as $O(1/\varepsilon)$. On the other hand, for $p=1-R+\varepsilon$, if a rate $R$ code is $(p,L)$ list decodable, the list size $L$ must be exponential in the code length $n$. The quantity 1R1-R is therefore referred to as the list-decoding capacity, to characterize the limiting fraction of errors that can be list-decoded as a function of the code rate. Informally, a code that is list-decodable up to radius $p=1-R-\varepsilon$ with list size $O_{\varepsilon}(1)$, or even list size $n^{O_{\varepsilon}(1)}$ where $n$ is the code length, is said to achieve (list-decoding) capacity.
@@
The list-decodability of Reed–Solomon codes is important for several reasons. Reed–Solomon codes are the most fundamental algebraic error-correcting codes. In fact, all of the prior explicit constructions of codes achieving list-decoding capacity are based on algebraic constructions that generalize Reed–Solomon codes, for example, Folded Reed–Solomon codes [34, 47], Multiplicity codes [38, 46, 47], and algebraic-geometric codes [15, 39, 40, 41]. Thus, it is natural to wonder whether and when Reed–Solomon codes themselves achieve list-decoding capacity. Additionally, all Reed–Solomon codes are optimally unique-decodable, so (equivalently) they are optimally list-decodable with the list size $L=1$, making them a natural candidate for codes achieving list-decoding capacity. Further, capacity-achieving Reed–Solomon codes would potentially offer advantages over existing explicit capacity-achieving codes, such as simplicity and potentially smaller alphabet sizes (which we achieve in this work). Lastly, list-decoding of Reed–Solomon codes has found several applications in complexity theory and pseudorandomness [10, 61, 50].
@@
For all these reasons, the list-decodability of Reed–Solomon codes is well-studied. As rate $R$ Reed–Solomon codes are uniquely decodable up to the optimal radius $\frac{1-R}{2}$ given by the Singleton Bound, the Johnson-bound [44] automatically implies that Reed–Solomon codes are $(p,L)$-list-decodable for error parameter $p=1-\sqrt{R}-\varepsilon$ and list size $L=O(1/\varepsilon)$. Guruswami and Sudan [36] showed how to efficiently list-decode Reed–Solomon codes up to the Johnson radius $1-\sqrt{R}$. For a long time, this remained the best list-decodability result (even non-constructively) for Reed–Solomon codes.
@@
Since then, several results suggested Reed–Solomon codes could not be list-decoded up to capacity, and in fact, not much beyond the Johnson radius $1-\sqrt{R}$. Guruswami and Rudra [33] showed that, for a generalization of list-decoding called list-recovery, Reed–Solomon codes are not list-recoverable beyond the (list-recovery) Johnson bound in some parameter settings. Cheng and Wan [12] showed that efficient list-decoding of Reed–Solomon codes beyond the Johnson radius in certain parameter settings implies fast algorithms for the discrete logarithm problem. Ben-Sasson, Kopparty, and Radhakrishnan [3] showed that full-length Reed–Solomon codes ($q=n$) are not list-decodable much beyond the Johnson bound in some parameter settings.
@@
Nevertheless, a subsequent exciting line of work [57, 59, 25, 19, 24, 9] has shown the existence of Reed–Solomon codes that could in fact be list-decoded beyond the Johnson radius. These works all consider combinatorial list-decodability of randomly punctured Reed–Solomon codes. By combinatorial list-decodability, we mean that the code is proved to be list-decodable without providing an algorithm to efficiently decode the list of nearby codewords. By randomly punctured Reed–Solomon code, we mean a code $\mathsf{RS}_{n,k}(\alpha_{1},\dots,\alpha_{n})$ where $(\alpha_{1},\dots,\alpha_{n})$ are chosen uniformly over all $n$-tuples of pairwise distinct elements of $\mathbb{F}_{q}$. Several of these works [57, 19, 24] proved more general list-decoding results about randomly puncturing any code with good unique-decoding properties, not just Reed–Solomon codes.
@@
In this line of work, a recent breakthrough of Brakensiek, Gopi, and Makam [9] showed, using notions of “higher-order MDS codes” [8, 56], that Reed–Solomon codes can actually be list-decoded up to capacity. In fact, they show, more strongly, that Reed–Solomon codes can be list-decoded with list size $L$ with radius $p=\frac{L}{L+1}(1-R)$, exactly meeting the generalized Singleton bound [59], resolving a conjecture of Shangguan and Tamo [59]. However, their results require randomly puncturing Reed–Solomon codes over an exponentially large field size $2^{O(n)}$, where $n$ is the block length of the code.
@@
A natural question is how small we can take the field size in a capacity-achieving Reed–Solomon code. It was shown [5, 1] that the exponential-in-$n$ field size in [9] is indeed necessary to exactly achieve the generalized Singleton bound [^2] In [5], this was shown for $L=2$ under the additional assumptions that the code is linear and MDS, and the general statement was later proved in [1]. but smaller field sizes remained possible if one allowed a small $\varepsilon$ slack in the parameters.
@@ 1.1 Our Results
List-decoding Reed–Solomon codes.
We show that Reed–Solomon codes are list-decodable up to capacity and the generalized Singleton bound with linear alphabet size $O(n)$, which is evidently optimal up to a constant factor. Our main result is the following.
@@
As in previous works like [9], Theorem 1.1 gives average-radius list-decodability, a stronger guarantee than list-decodability: for any distinct $L+1$ codewords $c^{(1)},\dots,c^{(L+1)}$ and any vector $y\in\mathbb{F}_{q}^{n}$, the average Hamming distance from $c^{(1)},\dots,c^{(L+1)}$ to $y$ is at least $\frac{L}{L+1}(1-R-\varepsilon)$. Taking $L=O(1/\epsilon)$ in Theorem 1.1, it follows that Reed–Solomon codes achieve list-decoding capacity even over linear-sized alphabets.
@@
Let $\varepsilon\in(0,1)$ and $q$ be a prime power such that $q\geq n+k\cdot 2^{O(1/\varepsilon^{2})}$. Then with probability at least $1-2^{-\Omega(n/\varepsilon)}$, a randomly punctured Reed–Solomon code of block length $n$ and rate $k/n$ over $\mathbb{F}_{q}$ is $\left(1-R-\varepsilon,O\left(\frac{1}{\varepsilon}\right)\right)$ average-radius list-decodable.
@@
In our proof of Theorem 1.1, we maintain a hypergraph perspective of the list-decoding problem, which was introduced in [25]. Section 2.2 elaborates on the advantages of this perspective, which include (i) more compact notations, definitions, and lemma statements, (ii) some more streamlined proofs, and (iii) an alternate presentation of ideas from Brakensiek, Gopi, and Makam [9] that more directly connects the list-decoding problem to the so-called GM-MDS theorem [13, 49, 65] via a hypergraph orientation theorem (see Appendix A).
List-decoding random linear codes.
@@
A random linear code of rate $R$ and length $n$ over $\mathbb{F}_{q}$ is a random subspace of $\mathbb{F}_{q}^{n}$ of dimension $Rn$. List-decoding random linear codes is well-studied [66, 17, 29, 28, 63, 57, 58, 48, 51, 31, 32, 53] and is an important question for several reasons. First, finding explicit codes approaching list-decoding capacity is a major challenge, and random linear codes provide a stepping stone towards explicit codes: it is easily seen that uniformly random codes achieve list-decoding capacity, and showing list-decodability of random linear codes can be viewed as a derandomization of the uniformly random construction (see [17] for a discussion of the challenge of showing list-decodability of linear codes, as was first done in [66]). Mathematically, the list-decodability of random linear codes concerns a fundamental geometric question: to what extent do random subspaces over $\mathbb{F}_{q}$ behave like uniformly random sets? In coding theory, list-decodable random linear codes are useful building blocks in other coding theory constructions [30, 42]. Lastly, the algorithmic question of decoding random linear codes is closely related to the Learning With Errors (LWE) problem in cryptography [55] and Learning Parity with Noise (LPN) problem in learning theory [4, 18].
@@
The list-decodability of random linear codes is more difficult to analyze than uniformly random codes, because codewords do not enjoy the same independence as in random codes. Thus the naive argument that shows that random linear codes achieve list-decoding capacity [66] gives an exponentially worse list size of $q^{1/\varepsilon}$ than for random codes ($\varepsilon$ is the gap to the “$q$-ary capacity”, $R=1-H_{q}(p)$, where $H_{q}(x)\stackrel{{\scriptstyle\rm def}}{{=}}x\log_{q}(q-1)-x\log_{q}(x)-(1-x)\log_{q}(1-x)$ is the $q$-ary entropy function). Several works have sought to circumvent this difficulty [17, 29, 28, 63, 57, 58, 48, 31] improving the list-size bound to $O_{q}(1/\varepsilon)$, matching the list-size of uniformly random codes.
@@
However, these results are more relevant for smaller alphabet sizes $q$, and approaching the alphabet-independent capacity of p=1Rp=1-R is less understood. In this setting, uniformly random codes are, with high probability, list-decodable to capacity with optimal alphabet size $2^{O(1/\varepsilon)}$  [^3] This follows from the list-decoding capacity theorem [35]. Over $q$-ary alphabets, the list-decoding capacity is given by $p=H_{q}^{-1}(1-R)$, which is larger than $1-R-\varepsilon$ when $q\geq 2^{\Omega(1/\varepsilon)}$. and optimal list size $O(1/\varepsilon)$. [^4] For codes over smaller alphabets, the list size $O(1/\varepsilon)$, where $\varepsilon$ is the gap to capacity, is believed to be optimal, but a proof is only known for large radius [37]. However, for approaching the alphabet-independent capacity, the list size $O(1/\varepsilon)$ is known to be optimal by the generalized Singleton bound [59]. However, it was not known whether random linear codes (or, in general, more structured codes) could achieve similar parameters. In particular, both of the following questions were open (as far as we are aware).
@@
  • Are rate $R$ random linear codes $(1-R-\varepsilon,O(1/\varepsilon))$-list-decodable with high probability? Previously, this was not known for any alphabet size $q$, even alphabet size growing with the length of the code. Previously, the best list size for random linear codes list-decodable to radius $p=1-R-\varepsilon$ was at least $2^{\Omega(1/\varepsilon)}$ [28, 58]. [^5] Prior works on list-decoding random linear codes were more relevant for $q$ an absolute constant such as $2,3,4,5,\dots$. [28] appears to give a list-size bound of $O(q^{O_{R}(1)}/\varepsilon)$, and [58] appears to give a list size bound that is at least $q^{\log^{2}(1/\varepsilon)}$, and we need $q\geq 2^{\Omega(1/\varepsilon)}$.
@@
By taking $L=O(1/\varepsilon)$, we see that random linear codes achieve capacity with optimal list size $O(1/\varepsilon)$ and near-optimal alphabet size $2^{O(1/\varepsilon^{2})}$.
@@
Alphabet size lower bounds.
Above, we saw that random linear codes achieve list-decoding capacity with optimal list-size and linear alphabet size. A natural question is determining the optimal constant in the alphabet size. We showed that $q\geq n\cdot 2^{O(1/\varepsilon^{2})}$ suffices, and by the list-decoding capacity theorem [17] — which requires $q\geq 2^{\Omega(1/\varepsilon)}$ — we cannot have better than an exponential-type dependence on $1/\varepsilon$ for subconstant $\varepsilon<O(1/\log n)$.
For approaching capacity with constant $\varepsilon$, Ben-Sasson, Kopparty, and Radhakrishnan [3] showed that, for any $c\geq 1$, there exist full-length Reed–Solomon codes that are not list-decodable much beyond the Johnson bound with list-sizes $O(n^{c})$. Thus in order to achieve list-decoding capacity, one needs $q>n$ in some cases. However, while full-length Reed–Solomon codes could not achieve capacity, perhaps it was possible that Reed–Solomon codes over field size, say $q=2n$ or even $q=(1+\gamma)n$, could achieve capacity in all parameter settings. We observe that, as a corollary of [3], such a strong guarantee is not possible: for any $c>1$, there exist a constant rate $R=R(c)>0$ and infinitely many field sizes $q$ such that all Reed–Solomon codes of length $n\geq q/c$ and rate $R$ over $\mathbb{F}_{q}$ are not list-decodable to capacity 1R1-R with list size $n^{c}$. The proof is in Section 5.
@@
Follow-up works.
This paper is based on two conference papers [27, 2]. The paper [27] proved Theorem 1.1 with a quadratic alphabet size, while [2] improved the alphabet size to linear and extended the techniques to random linear codes. There have already been several follow-ups to these works.
Brakensiek, Dhar, Gopi, and Zhang [7] proved that Algebraic Geometry (AG) codes achieve list-decoding capacity over constant-sized alphabets by combining our techniques with a generalized GM-MDS theorem, which Brakensiek, Dhar, and Gopi proved in [6].
Recently, Guo, Xing, Yuan, and Zhang [26] initiated the study of “higher-order MRD codes” as counterparts to higher-order MDS codes in the rank metric. They proved that random Gabidulin codes are list-decodable to capacity in the rank metric.